# ROSflight 2.0: Lean ROS 2-Based Autopilot for Unmanned Aerial Vehicles

### v1

@@ I Introduction
ROSflight is a lean, open-source autopilot for UAVs built for research. Because it is designed to be lean, ROSflight is not full-featured and does not boast many of the state-of-the-art functions available in other popular open-source autopilots [1, 2]. Instead, ROSflight offers only the basic functionality needed to support UAV research, prioritizing understandability. While this places more responsibility on the end user to develop application-specific code, we believe a lean architecture reduces the black-box nature of the autopilot, thus reducing the total effort to implement a user’s application code. Additionally, the ROSflight project has a strong emphasis on clear code and complete documentation, which improves accessibility and lowers the barrier to entry for researchers. Documentation can be found on the project website, rosflight.org.
@@
The rest of the paper is organized as follows: Section II describes work similar to ROSflight. A brief overview of ROSflight is described in Section III Improvements to the ROSflight architecture are described in detail in Section IV, and supported hardware and improvements to the simulation environment are discussed in Section V. Hardware results demonstrating these improvements to ROSflight are described in Section VI and we conclude in Section VII.
@@ II Related Work
Advances in UAV technology have given rise to many excellent and mature open-source autopilots like PX4 [1] and ArduPilot [2]. Both of these autopilots offer impressive feature sets, state-of-the-art performance, extensive community support and adoption, and plug-and-play functionality. Additionally, both projects have good support for the Robot Operating System (ROS 2), thus improving integration of external code libraries into an autopilot [6]. However, the large code bases, complexity, and steep learning curve of these autopilots decrease understandability, which increases the time and effort required to integrate external code and conduct simulation and hardware tests. ROSflight is primarily designed to be understandable with clean, lean code and complete documentation. Additionally, ROSflight moves most of the autopilot stack to a Linux-based companion computer which has more compute power and is easier to develop on than an embedded microcontroller. ROSflight also emphasizes true software-in-the-loop simulation [5]. Other open-source autopilots [7, 8] are less mature, have limited flexibility and ROS 2 support, or are designed primarily for hobbyist use. ROSflight is built around ROS 2 and is explicitly designed to fill researchers’ needs.
@@ III ROSflight Overview
Figure 1: Overview of the ROSflight archtecture.
@@ IV-A ROS 1 to ROS 2
Previous versions of ROSflight used ROS 1, which is no longer supported by the maintainers of ROS. ROSflight now uses ROS 2, and primarily supports long-term support (LTS) versions of ROS 2, which are ROS 2 Humble with Ubuntu 22.04 and ROS 2 Jazzy with Ubuntu 24.04. ROS 2 offers several improvements over the previous ROS 1 system, making it more robust and reliable [3]. The core ROSflight firmware on the embedded microcontroller does not use ROS (e.g. micro-ROS [10]), and instead communicates with the ROS 2-based I/O node (called ROSflightIO) on the companion computer over serial using MAVlink. The ROSflight simulation is built around ROS 2, and exploits this to improve modularity as will be discussed in Section V-B.
@@ IV-B Mixer
In ROSflight, we let $m=6$ and $n=10$, meaning there are ten output channels corresponding to hardware outputs (PWM or DShot) on the FCU, and six desired commands to the mixer, requiring 60 values. Additionally, the ROSflight mixer defines a PWM rate and an output type (motor, servo, GPIO, or auxiliary) for each output channel, yielding an additional 20 values. Note that the actual number of PWM output pins is limited by the FCU hardware. If a hardware board has more than ten PWM output pins, the remainder are automatically defined to be auxiliary outputs, which are simply PWM outputs that do not pass through the mixer.
As in [11], $u$ is typically a vector of the desired forces, $F$, and torques, $\tau$, that are to be applied to the physical system. Thus, in ROSflight, we let $u=[F^{T},\tau^{T}]^{T}\in\mathbb{R}^{6}$ as the default interpretation of the mixing equations. However, the actual meaning of the mixer input vector $u$ is closely tied with the elements of the mixer matrix. Thus, in the following descriptions, we do not constrain the input vector $u$ to be a vector of forces and torques, but instead a vector of six generic control inputs, all of which may not be used in a given setting. For example, a quadrotor mixer may take in desired forces and torques ($F_{z},Q_{x},Q_{y},Q_{z}$), corresponding to four of the six degrees of freedom of the physical aircraft, and return individual motor commands, while a V-tail fixed-wing mixer may take in desired control surface commands ($\delta_{a},\delta_{e},\delta_{t},\delta_{r}$) and return actual servo commands. An example of Equation (1) for a V-tail fixed-wing mixer is
$\begin{bmatrix}\delta_{a}\\ \delta_{lr}\\ \delta_{rr}\\ \delta_{t}\\ 0_{6\times 1}\end{bmatrix}=\begin{bmatrix}1&0&0&0&0_{1\times 2}\\ 0&-0.5&0.5&0&0_{1\times 2}\\ 0&0.5&0.5&0&0_{1\times 2}\\ 0&0&0&1&0_{1\times 2}\\ 0_{6\times 1}&0_{6\times 1}&0_{6\times 1}&0_{6\times 1}&0_{6\times 2}\end{bmatrix}\begin{bmatrix}\delta_{a}^{c}\\ \delta_{e}^{c}\\ \delta_{r}^{c}\\ \delta_{t}^{c}\\ 0_{2\times 1}\end{bmatrix}$,
where $u=[\delta_{a}^{c},\delta_{e}^{c},\delta_{r}^{c},\delta_{t}^{c},0_{1\times 2}]^{T}$ is the input vector containing the desired aileron, elevator, rudder, and throttle commands, respectively, and $\tau=[\delta_{a},\delta_{lr},\delta_{rr},\delta_{t},0_{1\times 6}]^{T}$ is the output vector containing the commands to the servos controlling the ailerons, left ruddervator, right ruddervator, and throttle, respectively. Since in this example the remaining two control inputs in $u$ are not used, they are zero. Note that the signs of the coefficients in the mixing matrix are dependent on the orientation of the control-surface servos on the aircraft.
@@ IV-B1 Predefined Mixers
Some predefined, hard-coded mixers are included for convenience in the implementation of the ROSflight firmware. These mixers can be selected at runtime via parameters, and are available for standard airframes, including quadorotor, hexarotor, standard fixed-wing, or V-tail fixed-wing airframes. A full list of predefined mixers is available on rosflight.org. Each mixer definition defines the type of each output channel (e.g., motor, servo, GPIO, or auxiliary), the PWM rate assigned to each channel, and the values of either $M$ or $M^{\dagger}$. See the appendix for the derivation and assumptions of the predefined mixers.
@@ IV-B2 Custom Mixer
Since the mixer is intimately tied to the physical geometry of the aircraft, users may need mixers not included in the predefined mixers hard-coded in ROSflight. Thus, ROSflight includes the ability for users to load their own custom mixing matrix values via parameters. A custom mixer must include both the 60 mixing matrix parameters as well as the 20 header parameters. The names of these parameters and default values can be found on rosflight.org.
@@ IV-B3 Pass-through Mixer
ROSflight can be operated in pass-through mode, where offboard commands bypass the firmware controller and progress directly to the mixer, as shown in Figure (2). The high-speed serial connection enables sending these pass-through offboard commands at a high rate, thus giving direct access to individual motors at a high rate. Such a configuration enables low-level control of a vehicle from the companion computer, thus improving flexibility and lowering the barrier to entry since all control loops can be closed on the Linux-based companion computer. Use cases might include sending direct actuator commands for individual motor control during transition of an eVTOL, or for sending actuator commands from the output of a neural network controller [12, 13]. Additionally, operating ROSflight in pass-through mode enables users to filter the inputs to the mixer. This could be used to simulate hardware failures [14], cyber attacks [15], or other physical failures simply by filtering the companion computer’s output before sending to the mixer.
@@ IV-B4 Primary and Secondary Mixers
When conducting flight tests with a companion computer, it can be helpful to isolate certain channels of the offboard command to incrementally test companion computer control. To accommodate this, the RC pilot in ROSflight has two types of control over the vehicle: attitude override and throttle override. When RC attitude override is enabled, the RC pilot has control over the $Q_{i}$ portions of the command vector $u$. When RC throttle override is enabled, the RC pilot has control over the $F_{i}$ portions of the command vector. When either override is enabled, the mixer that transforms the control input vector $u=[F_{x},F_{y},F_{z},Q_{x},Q_{y},Q_{z}]^{T}$ to motor outputs $\tau$ is a mixture of the primary and secondary mixers, as shown in Figure (3).
@@
This allows the RC pilot to control either the throttle or the attitude while the companion computer controls the other. Note that this requires that the mixing matrices and the control inputs are designed so that this split may occur. In other words, if the mixer output does not have meaning when the torque commands $Q_{i}$ are separated from the force commands $F_{i}$, then attempting to use attitude and throttle override independently of each other may cause undesirable behavior. If that is the case, setting the attitude and throttle override RC channels to the same value (thus activating both or neither) would lead to the correct interaction between mixers.
@@ IV-B5 Empirical Motor Parameters
As described in the appendix, using the motor and propeller parameters can increase the fidelity of the mixing matrix. In ROSflight, the USE_MOTOR_PARAM parameter tells the firmware whether or not the mixer directly outputs motor setpoints, $\delta_{t,i}$, or desired angular speeds, $\Omega_{i}$. If the mixer outputs desired angular speeds, then ROSflight converts these values to desired motor setpoints before they are sent to the motors and ESCs.
V Integration Updates
@@ V-B Software-in-the-Loop Simulation
The ROSflight simulation package is flexible and modular so that users can customize it to their needs by supporting their own simulator or by including custom functionality. ROSflight natively supports three simulators out of the box: HoloOcean [17], a photorealistic visualizer built on Unreal Engine 5, Gazebo [18], a popular open-source simulator for robotics, and a custom lightweight simulator that uses ROS 2 RViz for visualization.
The ROSflight simulation is split into a series of modules which are implemented as separate ROS 2 nodes and are described in Table III. This modularity enables users to define which modules are loaded at runtime. Since different visualizers come with different capabilities, this allows visualizers to be exchanged with minimal effort. For example, Gazebo performs dynamic propagation as part of the visualization engine, while the lightweight RViz visualizer only displays the 3D motion of the aircraft. Thus, when using Gazebo with ROSflight, the dynamics module is not launched while the rest of the ROSflight simulation modules remain unchanged.
@@ VI-A Serial Delay
To measure the approximate serial delay, we modified ROSflight firmware to send all received offboard commands back to the companion computer. The companion computer first packed an offboard command message, marked the timestamp, and sent it to ROSflight firmware. The firmware parsed the message as part of its normal routine and immediately sent it back. Thus, each offboard command was sent twice across the serial connection. The companion computer received the message and computed the round-trip time (RTT). Note that the ROSflight firmware continued to run all of its normal operations during this experiment, including collecting, packing, and sending sensor data, responding to heartbeat requests, filling parameter requests, etc. This means that some variability and delay is included in the RTT measurement, as well as the time required for the normal operation of ROSflight. Data was collected for 45 seconds, and the average, maximum, and minimum RTT values were recorded. Table IV shows RTT statistics for when the companion computer sends offboard commands at 400 Hz. A Rasberry Pi 5 was used as the companion computer for all configuration 1 tests.
@@
Figure 5 shows a histogram of the recorded RTT values for the data presented in Table IV, with commands being sent at 400 Hz. This shows that the maximum value of 56.921 ms reported in Table IV for configuration 1 is an outlier. Configuration 2 demonstrated significantly superior RTT values over configuration 1. Note that configuration 2 has an integrated design, but configuration 1 relies on a physical USB wire for the serial connection.
@@ VII Summary
This appendix contains the derivation of the general-form and predefined mixer equations for standard multirotor vehicles used in ROSflight. Also note that this derivation assumes that all outputs are motor outputs, not servo outputs. The fixed-wing mixer derivation is not included here, but the mixer equations for a general aircraft can be derived by following a similar procedure.
@@
Equation (8) shows the $i$th column of a simplified $M$, where the rows correspond to geometric properties associated with the given airframe. The assumptions made in the derivation of this equation limit the accuracy of this mixer model, but allow for an easy and intuitive understanding of the mixing matrix. The hardcoded mixing matrices (for multirotors) in ROSflight follow this form, allowing users to easily determine the mixing matrix and its limitations without knowledge of the motor parameters of the system.

### v2

@@ I Introduction
ROSflight is a lean, open-source autopilot for UAVs built for research. Because it is designed to be lean, ROSflight is not full-featured and does not boast many of the state-of-the-art functions available in other popular open-source autopilots [17, 2]. Instead, ROSflight offers only the basic functionality needed to support UAV research, prioritizing understandability. While this places more responsibility on the end user to develop application-specific code, we believe a lean architecture reduces the black-box nature of the autopilot, thus reducing the total effort to implement a user’s application code. Additionally, the ROSflight project has a strong emphasis on clear code and complete documentation, which improves accessibility and lowers the barrier to entry for researchers and students alike. Documentation can be found on the project website, rosflight.org.
@@
The rest of the paper is organized as follows: Section II describes work similar to ROSflight. A brief overview of ROSflight is described in Section III. Improvements to the ROSflight architecture are described in detail in Section IV, and supported hardware and improvements to the simulation environment are discussed in Section V. Hardware results demonstrating these improvements to ROSflight are described in Section VI and we conclude in Section VII.
@@ II Related Work
Advances in UAV technology have given rise to many excellent and mature open-source autopilots like PX4 [17] and ArduPilot [2]. Both of these autopilots offer impressive feature sets, state-of-the-art performance, extensive community support and adoption, and plug-and-play functionality. Additionally, both projects have good support for ROS 2, thus improving integration of external code libraries into an autopilot [1]. However, the large code bases, complexity, and steep learning curve of these autopilots decrease understandability. This can increase the time and effort required to integrate external code and conduct simulation and hardware tests. As noted by D’Angelo et al., these challenges often cause researchers to prefer developing custom firmware from scratch [5].
ROSflight is primarily designed to be understandable with clean, lean code and complete documentation. Since ROSflight is lean, it does not offer most of the features offered by [17, 2], and therefore is not suitable for commercial applications or production-ready vehicles. Instead, ROSflight targets research, classroom, or early-stage development projects where understandability and ease of modification is crucial.
Additionally, ROSflight moves most of the autopilot stack to a Linux-based companion computer which has more compute power and is easier to develop and rapidly test than an embedded microcontroller. ROSflight also emphasizes true software-in-the-loop simulation [11], meaning that the core software that runs in simulation also runs in hardware, with no changes. Other open-source autopilots [20, 9] are less mature, have limited flexibility and ROS 2 support, or are designed primarily for hobbyist use. ROSflight is built around ROS 2 and is explicitly designed to fill researchers’ needs.
@@ III ROSflight Overview
Figure 1: Overview of the ROSflight architecture.
@@ IV-A ROS 1 to ROS 2
Previous versions of ROSflight used ROS 1, which is no longer supported by the maintainers of ROS. ROSflight now uses ROS 2, and primarily supports long-term support (LTS) versions of ROS 2, which are ROS 2 Humble with Ubuntu 22.04 and ROS 2 Jazzy with Ubuntu 24.04. ROS 2 offers several improvements over the previous ROS 1 system, making it more robust and reliable [15]. The core ROSflight firmware on the embedded microcontroller does not use ROS (e.g. micro-ROS [4]), and instead communicates with the ROS 2-based I/O node (called ROSflightIO) on the companion computer over a serial connection. The ROSflight simulation is built around ROS 2, and exploits this to improve modularity as will be discussed in Section V-B.
@@ IV-B Mixer
In ROSflight, we let $m=10$ and $n=10$, meaning there are ten output channels corresponding to hardware outputs (PWM or DShot) on the FCU, and ten desired commands to the mixer, requiring 100 values. Additionally, the ROSflight mixer defines a PWM rate and an output type (motor, servo, GPIO, or auxiliary) for each output channel, yielding an additional 20 values. Note that the actual number of PWM output pins is limited by the FCU hardware. If a hardware board has more than ten PWM output pins, the remainder are automatically defined to be auxiliary outputs, which are simply PWM outputs that do not pass through the mixer.
The authors in [3] define $u$ to be a vector of the desired forces, $F$, and torques, $Q$, that are to be applied to the physical system. However, the semantic meaning of the mixer input vector $u$ is closely tied with the elements of the mixer matrix. Thus, ROSflight does not constrain the input vector $u$ to be a vector of forces and torques, but instead a vector of ten generic control inputs, all of which may not be used in a given setting. For example, a quadrotor mixer may take in desired forces and torques ($F_{z},Q_{x},Q_{y},Q_{z}$), corresponding to four of the six degrees of freedom of the physical aircraft, and return individual motor commands, while a V-tail fixed-wing mixer may take in desired control surface commands ($\delta_{a},\delta_{e},\delta_{t},\delta_{r}$) and return individual servo commands. An example of Equation (1) for a V-tail fixed-wing mixer is
$\begin{bmatrix}\delta_{a}\\ \delta_{lr}\\ \delta_{rr}\\ \delta_{t}\\ 0_{6\times 1}\end{bmatrix}=\begin{bmatrix}1&0&0&0&0_{1\times 6}\\ 0&-0.5&0.5&0&0_{1\times 6}\\ 0&0.5&0.5&0&0_{1\times 6}\\ 0&0&0&1&0_{1\times 6}\\ 0_{6\times 1}&0_{6\times 1}&0_{6\times 1}&0_{6\times 1}&0_{6\times 6}\end{bmatrix}\begin{bmatrix}\delta_{a}^{c}\\ \delta_{e}^{c}\\ \delta_{r}^{c}\\ \delta_{t}^{c}\\ 0_{6\times 1}\end{bmatrix}$,
where $u=[\delta_{a}^{c},\delta_{e}^{c},\delta_{r}^{c},\delta_{t}^{c},0_{1\times 6}]^{T}$ is the input vector containing the desired aileron, elevator, rudder, and throttle commands, respectively, and $\tau=[\delta_{a},\delta_{lr},\delta_{rr},\delta_{t},0_{1\times 6}]^{T}$ is the output vector containing the commands to the servos controlling the ailerons, left ruddervator, right ruddervator, and throttle, respectively. Since in this example the remaining six control inputs in $u$ are not used, they are zero. Note that the signs of the coefficients in the mixing matrix are dependent on the orientation of the control-surface servos on the aircraft.
The command vector $u$ is not constrained to be an ordered vector of forces and torques, so ROSflight must determine how to map RC inputs to the $u$ command vector. To do this, ROSflight assumes that the RC command vector takes the form $u=[F_{x},F_{y},F_{z},Q_{x},Q_{y},Q_{z},0_{1\times 4}]^{T}$, where the $F$ channels correspond to force or throttle RC inputs. The $Q$ channels correspond to torque or rotational RC inputs (e.g. roll torque, angle, or rate commands).
@@ IV-B1 Predefined Mixers
Some predefined, hard-coded mixers are included for convenience in the implementation of the ROSflight firmware. These mixers can be selected at runtime via parameters, and are available for standard airframes, including quadrotor, hexarotor, standard fixed-wing, or V-tail fixed-wing airframes. A full list of predefined mixers is available on rosflight.org. Each predefined mixer definition defines the type of each output channel (e.g., motor, servo, GPIO, or auxiliary), the PWM rate assigned to each channel, and the values of either $M$ or $M^{\dagger}$. See [18] for the derivation and assumptions of the predefined mixers.
@@ IV-B2 Custom Mixer
Since the mixer is intimately tied to the physical geometry of the aircraft, users may need mixers not included in the predefined mixers hard-coded in ROSflight. Thus, ROSflight includes the ability for users to load their own custom mixing matrix values via parameters. A custom mixer must include both the 100 mixing matrix parameters as well as the 20 header parameters. The names of these parameters and default values can be found on rosflight.org.
@@ IV-B3 Pass-through Mixer
ROSflight can be operated in pass-through mode, where offboard commands bypass the firmware controller and progress directly to the mixer, as shown in Figure (2). The high-speed serial connection enables sending these pass-through offboard commands at a high rate, thus giving direct access to individual motors at a high rate. Such a configuration enables low-level control of a vehicle from the companion computer. Control over this low-level behavior is advantageous when developing novel flight controllers; however, this is not always possible with state-of-the-art autopilots like PX4 without significant changes to the firmware and navigating complex dependencies [8]. Additionally, PX4 explicitly disallows pass-through control in some circumstances [22].
ROSflight avoids this complexity by allowing all control loops to be closed on the Linux-based companion computer, thus improving flexibility. Use cases might include sending direct actuator commands for individual motor control during transition of an eVTOL, or for sending actuator commands from the output of a neural network controller [12, 14]. Additionally, operating ROSflight in pass-through mode enables users to filter the inputs to the mixer. This could be used to simulate hardware failures [19], cyber attacks [16], or other physical failures simply by filtering the companion computer’s output before sending to the mixer.
@@ IV-B4 Primary and Secondary Mixers
When conducting flight tests with a companion computer, it can be helpful to isolate certain channels of the offboard command to incrementally test companion computer control. To accommodate this, the RC pilot in ROSflight has two types of control over the vehicle: attitude override and throttle override. When RC attitude override is enabled, the RC pilot has control over the attitude ($Q$) portions of the command vector $u$. When RC throttle override is enabled, the RC pilot has control over the throttle ($F$) portions of the command vector. When either override is enabled, the mixer that transforms the control input vector $u=[F_{x},F_{y},F_{z},Q_{x},Q_{y},Q_{z},0_{1\times 4}]^{T}$ to motor outputs $\tau$ is a mixture of the primary and secondary mixers, as shown in Figure (3).
@@
This allows the RC pilot to control either the throttle or the attitude while the companion computer controls the other. Note that this requires that the mixing matrices and the control inputs are designed so that this split may occur. In other words, if the mixer does not have proper meaning as a combination of the primary and secondary mixers, then attempting to use attitude and throttle override independently of each other may cause undesirable behavior. If that is the case, setting the attitude and throttle override RC channels to the same value (thus activating both or neither) would result in either the entire primary mixer or the entire secondary mixer being used.
@@ IV-B5 Empirical Motor Parameters
As described in [18], using the motor and propeller parameters can increase the fidelity of the mixing matrix. In ROSflight, the USE_MOTOR_PARAM parameter tells the firmware whether or not the mixer directly outputs motor setpoints, $\delta_{t,i}$, or desired angular speeds, $\Omega_{i}$. If the mixer outputs desired angular speeds, then ROSflight converts these values to desired motor setpoints before they are sent to the electronic speed controllers (ESC).
V Hardware Integration
@@ V-B Software-in-the-Loop Simulation
ROSflight natively supports three simulators out of the box: HoloOcean [21], a photorealistic visualizer built on Unreal Engine 5, Gazebo [13], a popular open-source simulator for robotics, and a custom lightweight simulator that uses ROS 2 RViz for visualization. Figure (5) shows each of these three supported environments.
(a) Multirotor flying a waypoint mission in the RViz simulation environment.
(b) Fixed-wing aircraft flying a waypoint mission in the Gazebo simulation environment.
(c) Multirotor flying in the photorealistic Unreal Engine 5-based HoloOcean simulation environment. Camera images can be easily simulated.
Figure 5: Simulation environments supported natively by ROSflight, from photorealistic to lightweight environments that can be run headlessly.
Since different projects often require simulation environments with different features, the ROSflight simulation package is flexible and modular so that users can support their own simulator or include custom functionality. To do this, the ROSflight simulation is split into a series of modules which are implemented as separate ROS 2 nodes and are described in Table III. Figure (6) shows how each module interacts and what information is passed between modules. This modularity enables users to define which modules are loaded at runtime. Since different visualizers come with different capabilities, this allows visualizers to be exchanged with minimal effort. For example, Gazebo performs dynamic propagation as part of the visualization engine, while the lightweight RViz visualizer only displays the 3D motion of the aircraft. Thus, when using Gazebo with ROSflight, the dynamics module is not launched while the rest of the ROSflight simulation modules remain unchanged.
Figure 6: Diagram of the modules of the ROSflight simulation. Each module has a single responsibility, enabling easier modification by reducing dependencies between modules.
@@ VI-A Serial Delay
To measure the approximate serial delay, we modified ROSflight firmware to send all received offboard commands back to the companion computer. The companion computer first packed an offboard command message, marked the timestamp, and sent it to ROSflight firmware. The firmware parsed the message as part of its normal routine and immediately sent it back. Thus, each offboard command was sent twice across the serial connection. The companion computer received the message and computed the round-trip time (RTT). Note that the ROSflight firmware continued to run all of its normal operations during this experiment, including collecting, packing, and sending sensor data, responding to heartbeat requests, filling parameter requests, etc. This means that some variability and delay is included in the RTT measurement, as well as the time required for the normal operation of ROSflight. Data was collected for 45 seconds, and the average, maximum, and minimum RTT values were recorded. Table IV shows RTT statistics for when the companion computer sends offboard commands at 400 Hz. A Raspberry Pi 5 was used as the companion computer for all configuration 1 tests.
@@
Figure 7 shows a histogram of the recorded RTT values for the data presented in Table IV, with commands being sent at 400 Hz. This shows that the maximum value of 56.921 ms reported in Table IV for configuration 1 is an outlier. Configuration 2 demonstrated significantly superior RTT values over configuration 1. Note that configuration 2 has an integrated design, but configuration 1 relies on a physical USB connection for the serial connection.
@@ VII Summary
This appendix contains the derivation of the general-form and predefined mixer equations for standard multirotor vehicles used in ROSflight. This derivation assumes that all outputs are motor outputs, not servo outputs. The fixed-wing mixer derivation is not included here, but the mixer equations for a general aircraft can be derived by following a similar procedure.
@@
Equation (8) shows the $i$th column of a simplified $M$, where the rows correspond to geometric properties associated with the given airframe. The assumptions made in the derivation of this equation limit the accuracy of this mixer model, but allow for an easy and intuitive understanding of the mixing matrix. The hard-coded mixing matrices (for multirotors) in ROSflight follow this form, allowing users to easily determine the mixing matrix and its limitations without knowledge of the motor parameters of the system.